OpenStudy (anonymous):

A ball dropped from a height of 10 feet bounces back 9/10 of that distance. With each successive bounce, the ball continues to reach 9/10 of its previous height. What is the total vertical distance (both up and down) traveled by the ball when it stops bouncing? (Hint: Add the total distance the ball falls to the total distance it rises.)

OpenStudy (raden):

Total distance = 10 * (10+9)/(10-9) = 10 * 19 = 190 ft

OpenStudy (anonymous):

thank you

OpenStudy (raden):

you're welcome :)

OpenStudy (raden):

read this too to understanding your question above http://openstudy.com/users/raden#/updates/51888eabe4b0f910f30a8cc8